OpenStudy (anonymous):

19. Substitution would be the simplest method of solution for which system? (Points : 2) 5/3 x = 6y and 2x -3y = 8 y = 2x and x + y = 28 3x - 2y =18 and 4x + 5y =60 20. The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? (Points : 2) 56 41 69 55

OpenStudy (rogue):

19) The second system.

OpenStudy (rogue):

20) Solve it like a system x + y = 110 x - y = 28 x = y + 28 ----> x + y = 110 y + 28 + y = 110 2y + 28 = 110 2y = 82 y = 41 x+ y = 110 x + 41 = 110 x = 69 Choice 2) 41

OpenStudy (anonymous):

FOR 2nd question: x+y=110, x-y=28, by adding them: 2x=138 x=138/2=69 y=110-x=110-69=41 x=69, y=41